For a particular peculiar pair of dice, the probabilities of rolling 1, 2, 3, 4, 5 and 6 on each die are in the ratio 1:2:3:4:5:6. What is the probability of rolling a total of 7 on the two dice?
Answer by edjones(8007)

6 1 and 1 6 : 1/21 * 6/21 * 2 = .0272
5 2 and 2 5 : 2/21 * 5/21 * 2 = .0454
4 3 and 3 4 : 3/21 * 4/21 * 2 = .0544
.0272+.0454+.0544=.1270
.
Ed